# Uplift modeling with continuous treatments: A predict-then-optimize approach

### v1

@@ 1 Introduction
In many applications, decision-makers are interested in learning the causal effects of a treatment on a particular outcome (Holland, 1986; Pearl, 2009). A crucial aspect is response heterogeneity where entities respond differently to treatments based on their characteristics (Rubin, 2005). In intervention-based decision-making, precisely this heterogeneity is valuable to leverage for finding a treatment assignment policy that optimizes a specific outcome variable, based on historical data, while adhering to certain constraints such as scarce resources. Uplift modeling (UM) is a set of techniques to find such policies (Devriendt et al., 2018). An often-used predict-then-optimize UM approach combines conditional average treatment effect (CATE) estimation with an optimization step, where CATE estimation can be considered an inference step as part of a larger decision-making process (Fernández-Loría and Provost, 2022).
Many real-world applications, as illustrated by Table 1, contain complexities beyond binary treatments and budget as the sole constraint and are better characterized by a treatment dose, i.e., where the intervention can be applied across a continuous range of values (Hirano and Imbens, 2004; Holland-Letz and Kopp-Schneider, 2015). Moreover, continuous treatments have benefits for making treatment allocation possibly more effective and efficient compared to its binary counterpart because the marginal utility of treatment can vary with different dose levels and the treatments can be allocated on fine granularity.
@@
While there is a growing body of literature focusing on conditional average dose response (CADR) estimation (i.e., the continuous-valued counterpart of CATE), it focuses on causal inference and finding optimal doses, largely ignoring the role of continuous treatments in constrained decision-making contexts (Bica et al., 2020; Schwab et al., 2020; Nie et al., 2021). Constraints play a vital role in embedding business requirements into decision-making, which we accomplish using an integer linear programming (ILP) formulation. While many requirements exist, fairness stands out as a key consideration and is the main exemplary constraint throughout this work (Barocas et al., 2023). Though fairness constraints are not new in machine learning for decision-making, they are often embedded as soft constraints during model training (Han et al., 2023; Frauen et al., 2023). This approach, however, is limited in flexibility and modularity. By shifting the inclusion of fairness considerations to a post-processing step, i.e., after model training, decision-making is more flexible toward dynamic business requirements.
@@
    We provide a framework that clearly defines UM, and how this differentiates from a mere causal inference problem, i.e., treatment effect estimation;
@@
    We extend UM methods to effectively handle continuous treatments by first leveraging state-of-the-art causal ML methods for CADR estimation and then defining the optimization part as a dose-allocation problem. The ILP formulation offers flexibility to adapt the objective function (e.g., to incorporate cost-sensitivity) and to include additional constraints (e.g., to enforce fairness considerations);
@@
    We are the first to include fairness considerations in a UM setting as explicit constraints in an ILP formulation. By extension, to the best of our knowledge, this is also the first application of such constraints in decision-making pipelines that utilize ML predictions as input;
@@
The remainder of this paper is structured as follows. Section 2 defines UM and elaborates on related literature. Section 3 introduces the problem formulation and establishes the necessary notation. Section 4 focuses on the methodology, explaining the predict-then-optimize approach, the predictive models used to estimate CADRs, and the optimization techniques employed for constrained treatment allocation. Section 5 details the experimental setup, including data, evaluation metrics, and the experiments, followed by a discussion of the results. Finally, Section 6 discusses our findings, their implications, and potential future research and concludes the paper.
@@ 2.1 Purpose and definition
UM is a well-established set of methods in the field of personalized decision-making, closely aligned with the goals of prescriptive analytics (Devriendt et al., 2018). Unlike traditional predictive models that identify entities likely to yield a desired outcome, UM distinguishes between baseline responders (entities likely to show positive outcomes even without intervention) and true responders that respond because of the treatment. Consequently, UM focuses on outcome changes directly attributable to the treatment assignment, rather than simply predicting positive outcome probabilities. This helps avoid suboptimal targeting and ensures assignment policies capture their real incremental impact. Traditional UM predominantly deals with binary treatments, where the main goal is to rank individuals based on their expected response to a treatment (Zhang et al., 2021). Works such as Devriendt et al. (2018) provide a comprehensive overview, focusing on a setting with binary treatment effects, various modeling techniques, and the associated challenges. However, the literature often lacks a unified definition of UM, with some studies equating it to treatment effect inference (Gutierrez and Gérardy, 2017) while others focus primarily on ranking methods (Devriendt et al., 2020). To resolve this ambiguity, we propose the following UM definition:
@@
Uplift modeling refers to the collection of methods where the task at hand is optimally allocating treatments, with the objective of maximizing the total benefit generated by these treatments, determined by the cumulative uplift over entities, within a given set of constraints.
The above definition encompasses both the one-step approach (i.e., predict-and-optimize) and the two-step approach (i.e., predict-then-optimize). As Table 2 highlights, the methodological contribution of this work focuses on the two-step approach, where treatment effect estimation is a component of the broader UM framework, which also includes treatment allocation as a critical task.
@@ 2.2 Treatment effects
A treatment refers to an intervention or action that can be applied to an entity to influence a particular outcome. Treatments can be binary, where the action is treat or do not treat (e.g., offering or withholding discount). The CATE measures the expected difference in potential outcomes between treated and untreated groups, conditioned on entities’ characteristics (Shalit et al., 2017). Additionally, recent developments have extended traditional UM to multi-treatment scenarios, which generalize the setup by considering more than two treatment options (Olaya et al., 2020) (e.g. offering discounts through different channels), and sequential treatments, where multiple treatments can be applied over time, each potentially influencing the outcome effects of subsequent treatments (Hatt and Feuerriegel, 2024). However, these are not the focus of our paper and are considered future extensions.
@@
The estimation of causal treatment effects from data is inherently challenging due to the fundamental problem of causal inference, namely, the absence of counterfactual outcomes (Holland, 1986). In the case of observational data, this problem is exacerbated by the possible non-random assignment of treatments and the resulting confounding bias (Pearl, 2009). A real-life example is the overestimation of workplace wellness programs’ impact on employee wellbeing, where non-random incentives cause treated populations to differ from the general population (Jones et al., 2019). A naive approach may overfit on the self-selected group, missing the true causal effect. This highlights how confounding can lead to issues like Simpson’s paradox, where aggregated data misrepresents trends compared to stratified data (Jones et al., 2019). While randomized controlled trials (RCTs) are the gold standard for mitigating such biases (Verstraete et al., 2023), they are often impractical in real-world business settings because of cost, ethical, or operational reasons. Moreover, in the context of continuous treatments, with virtually infinite possible doses, setting up an RCT for dose-response estimation becomes significantly more complex (Holland-Letz and Kopp-Schneider, 2015). In contrast, observational data is often cheap and readily accessible. Therefore, methods that aim to balance the treatment and control groups have been proposed with techniques like propensity score matching (Rosenbaum and Rubin, 1983) or covariate balancing (Johansson et al., 2016; Shalit et al., 2017).
@@
In recent years, the consideration of continuous treatments—and consequently, CADR estimation—has gained traction because of its relevance in applications where understanding the effects of varying dose levels is critical (see Table 1). Estimating CADRs is inherently more challenging than CATEs because it involves a spectrum of treatment doses, rather than values 00 and $1$ (Schwab et al., 2020). Also, with the theoretically infinite number of possible dose values, data sparsity and the non-uniformity of dose assignment could be challenging, making accurate effect estimation at each dose level difficult (Bockel-Rickermann et al., 2024). The Conditional Average Dose Effect (CADE), which can be directly derived from the CADR, generalizes the CATE for continuous treatments. Our paper specifically focuses on UM with this type of treatment. Traditionally, dose effects have been derived from average dose-response curves using methods like the Hirano-Imbens estimator (Hirano and Imbens, 2004), which extends propensity scores to continuous interventions through generalized propensity scores. However, these approaches often overlook individual-level heterogeneity in responses. Recent advancements in ML have introduced techniques for learning individualized dose-response curves (Bica et al., 2020; Schwab et al., 2020; Nie et al., 2021; Zhang et al., 2022; Bockel-Rickermann et al., 2023). However, analogously to the case of CATE estimation, these methods primarily focus on the inference step and do not incorporate an optimization step where the learned dose-response relationships are used to prescribe decisions (Fernández-Loría and Provost, 2022). Moreover, unlike some prior work like Zhou et al. (2023), we do not assume that an increased dose always leads to greater benefits, and instead dose-response curves can be non-monotonic and vary in shape.
@@ 2.3.1 Cumulative uplift & total benefit as driver
The allocation task aims to optimize an objective value driven by (cumulative) uplift. While uplift applies to a single entity, cumulative uplift reflects the policy’s overall impact. For a policy treating $K$ entities, cumulative uplift is the sum of CATEs for discrete treatments or CADEs for continuous treatments across these treated entities. In certain applications, cumulative uplift directly aligns with the target objective, as seen in contexts where maximizing cumulative uplift is the explicit goal (Devriendt et al., 2018). However, cumulative uplift itself does not necessarily represent the objective, for example in value-driven analytics where uplift might disregard expected profit and in cases where each entity may be associated with an individual benefit and treatment cost (Gubela et al., 2020; Lemmens and Gupta, 2020; Gubela and Lessmann, 2021; Verbeke et al., 2023). Hence, when cumulative uplift is not the end goal in itself, allocations with lower cumulative uplift may yield a greater increase in objective value.
@@ 2.3.2 Predict-and-optimize
Although DFL approaches for UM exist (Betlei et al., 2021; Vanderschueren et al., 2024), we adopt the predict-then-optimize framework due to its simplicity, flexibility, stability, and generalizability to other treatment types. It allows using various predictive models without being restricted by a specific optimization setup, which is especially useful for complex or pre-trained models without access to the training process (Barocas et al., 2023). Moreover, it enables flexible adjustments to the objective function and integration of additional constraints other than budget limits, such as fairness considerations, allowing for different allocations under varying budget conditions without requiring a complete model retraining.
@@ 2.3.4 Flexible optimization constraints and objectives
Balancing utility—such as profit and accuracy—against fairness, like equal treatment across demographics, remains a key challenge. Kleinberg et al. (2016) and Corbett-Davies et al. (2017) show that enhancing fairness often trades off with utility and that multiple fairness definitions may even conflict. Traditional fairness assessments emphasize output-based metrics, such as demographic parity or equal opportunity, which measure group-level outcome disparities (Barocas et al., 2023). Recent approaches consider the entire distribution of predictions or decisions, incorporating statistical moments and exploring distributional fairness (Khan et al., 2023; Han et al., 2023). In decision-making, fairness extends beyond prediction to both allocation fairness—ensuring treatment levels are independent of sensitive attributes like race or gender—and outcome fairness, which ensures equitable results across groups. Fairness notions, which may conflict, can be incorporated as hard constraints during optimization (Corbett-Davies et al., 2017), similar to this work’s approach, or as soft constraints in multi-objective learning (Barocas et al., 2023). Recent research, including Frauen et al. (2023), has largely focused on fairness in observational data with binary sensitive attributes, while Nabi et al. (2019) apply causal inference to multi-stage decision-making under fairness constraints. However, these studies do not address fairness in the context of continuous treatments, a critical gap our research aims to fill.
@@
In managerial decision-making, profit maximization or cost reduction are often the primary goals. Cost-sensitive or value-driven methods are crucial as they incorporate asymmetric costs and benefits, aligning decisions with these business objectives (Höppner et al., 2022). Cost-sensitive approaches seek to balance costs and benefits, an aspect often ignored by standard methods (Gubela et al., 2020; Verbeke et al., 2023). In this work, costs and benefits are deterministic while assuming stochastic treatment effects, setting it apart from studies where they are both modeled as stochastic (Verbraken et al., 2012; Verbeke et al., 2012). Cost-sensitive methods find applications across various domains, including credit risk (Bahnsen et al., 2014), fraud detection (Vasquez et al., 2022; De Vos et al., 2023), customer churn (Verbraken et al., 2012), business failure prediction (De Bock et al., 2020), and machine maintenance (Vanderschueren et al., 2023b).
@@ 3.2 Prediction step
During the prediction step, our goal is to train a model that can accurately estimate $\hat{y}(s)$, providing unbiased estimates of $\mu(s,\mathbf{x})$ and $\tau_{s}(\mathbf{x})$ across the entire domain of $S\in[0,1]$. The estimated CADR for an instance with features $\mathbf{x}$ is defined as $\hat{\mu}(s,\mathbf{x})=\mathbb{E}\left[\hat{Y}(s)\mid\mathbf{X}=\mathbf{x}\right]$. The estimated CADE for a given dose $s$ is defined as $\hat{\tau}_{s}(\mathbf{x})=\hat{\mu}(s,\mathbf{x})-\hat{\mu}(0,\mathbf{x})=\mathbb{E}\left[\hat{Y}(s)-\hat{Y}(0)\mid\mathbf{X}=\mathbf{x}\right]$. To use the CADE estimates as input for the optimization step, we discretize the CADRs into $\delta$ bins. For a given $\delta$, we define $D=\left\{\left.\frac{d-1}{\delta}\right\rvert\,d=1,\ldots,(\delta+1)\right\}$ so that the vector $\hat{\bm{\tau}}(\mathbf{x})=\Bigl{(}\hat{\tau}_{D_{d}}(\mathbf{x})\Bigr{)}_{d=1}^{\delta}$ then contains an entity’s $\delta$ CADE estimates.

### v2

@@ 1 Introduction
In many applications, decision-makers are interested in learning the causal effects of a treatment on a particular outcome (Holland, 1986; Pearl, 2009). A crucial aspect is response heterogeneity, where entities respond differently to treatments based on their characteristics (Rubin, 2005). In intervention-based decision-making, precisely this heterogeneity is valuable to leverage for finding a treatment assignment policy that optimizes a specific outcome variable, based on historical data, while adhering to certain constraints such as scarce resources. Uplift modeling (UM) is a set of techniques to find such policies (Devriendt et al., 2018). An often-used predict-then-optimize UM approach combines conditional average treatment effect (CATE) estimation with an optimization step, where CATE estimation can be considered an inference step as part of a larger decision-making process (Fernández-Loría and Provost, 2022).
Many real-world applications, as illustrated by Table 1, contain complexities beyond binary treatments and budget as the sole constraint and are better characterized by a treatment dose, i.e., where the intervention can be applied across a continuous range of values (Hirano and Imbens, 2004; Holland-Letz and Kopp-Schneider, 2015). Moreover, continuous treatments have benefits for making treatment allocation possibly more effective and efficient compared to their binary counterpart because the marginal utility of treatment can vary with different dose levels, and the treatments can be allocated on fine granularity.
@@
While there is a growing body of literature focusing on conditional average dose response (CADR) estimation (i.e., the continuous-valued counterpart of CATE), this line of work focuses on causal inference and finding optimal doses, largely ignoring the role of continuous treatments in constrained decision-making contexts (Bica et al., 2020; Schwab et al., 2020; Nie et al., 2021). Constraints play a vital role in embedding business requirements into decision-making, which we accomplish using an integer linear programming (ILP) formulation. While many requirements exist, fairness stands out as a key consideration and is the main exemplary constraint throughout this work (Barocas et al., 2023). Though fairness constraints are not new in machine learning for decision-making, they are often embedded as soft constraints during model training (Han et al., 2023; Frauen et al., 2023). This approach, however, is limited in flexibility and modularity. By shifting the inclusion of fairness considerations to a post-processing step, i.e., after model training, decision-making is more flexible toward dynamic business requirements.
@@
    We provide a framework that clearly defines UM, and how this differentiates from a mere causal inference problem, i.e., treatment effect estimation.
@@
    We extend UM methods to effectively handle continuous treatments by first leveraging state-of-the-art causal ML methods for CADR estimation and then defining the optimization part as a dose-allocation problem. The ILP formulation offers flexibility to adapt the objective function (e.g., to incorporate cost-sensitivity) and to include additional constraints (e.g., to enforce fairness considerations).
@@
    We are the first to include fairness considerations in a UM setting as explicit constraints in an ILP formulation. By extension, to the best of our knowledge, this is also the first application of such constraints in decision-making pipelines that utilize ML predictions as input.
@@
The remainder of this paper is structured as follows. Section 2 defines UM and elaborates on related literature. Section 3 introduces the problem formulation and establishes the necessary notation. Section 4 focuses on the methodology, explaining the predict-then-optimize approach, the predictive models used to estimate CADRs, and the optimization techniques employed for constrained treatment allocation. Section 5 details the experimental setup, including data, evaluation metrics, and the experiments, followed by a discussion of the results. Finally, Section 6 presents our conclusions, discusses limitations, and outlines potential further research.
@@ 2.1 Purpose and definition
UM is a well-established set of methods in the field of personalized decision-making, closely aligned with the goals of prescriptive analytics (Devriendt et al., 2018). Unlike traditional predictive models that identify entities likely to yield a desired outcome, UM distinguishes between baseline responders (entities likely to show positive outcomes even without intervention) and true responders that respond because of the treatment. Consequently, UM focuses on outcome changes directly attributable to the treatment assignment, rather than simply predicting positive outcome probabilities. This helps avoid suboptimal targeting and ensures assignment policies capture their true incremental impact. Traditional UM predominantly deals with binary treatments, where the main goal is to rank individuals based on their expected response to a treatment (Zhang et al., 2021). Works such as Devriendt et al. (2018) provide a comprehensive overview, focusing on a setting with binary treatment effects, various modeling techniques, and the associated challenges. However, the literature often lacks a unified definition of UM, with some studies equating it to treatment effect inference (Gutierrez and Gérardy, 2017) while others focus primarily on ranking methods (Devriendt et al., 2020). To resolve this ambiguity, we propose the following UM definition:
@@
Uplift modeling refers to the collection of methods where the task at hand is optimally allocating treatments, with the objective of maximizing the total benefit generated by these treatments, determined by the cumulative uplift over entities, under given constraints.
The above definition encompasses both the one-step approach (i.e., predict-and-optimize) and the two-step approach (i.e., predict-then-optimize). Notably, this definition is method-agnostic to keep it inclusive of decision-focused methods that learn policies directly and therefore do not output an explicit treatment-effect estimate. As Table 2 highlights, the methodological contribution of this work focuses on the two-step approach, where treatment effect estimation is a component of the broader UM framework, which also includes treatment allocation as a critical task.
@@ 2.2 Treatment effects
A treatment refers to an intervention or action that can be applied to an entity to influence a particular outcome. Treatments can be binary, where the action is treat or do not treat (e.g., offering or withholding discount). The CATE measures the expected difference in potential outcomes between treated and untreated groups, conditioned on entities’ characteristics (Shalit et al., 2017). Additionally, recent developments have extended traditional UM to multi-treatment scenarios, which generalize the setup by considering more than two treatment options (Olaya et al., 2020) (e.g., offering discounts through different channels), and sequential treatments, where multiple treatments can be applied over time, each potentially influencing the outcome effects of subsequent treatments (Hatt and Feuerriegel, 2024). However, these are not the focus of our paper and are considered future extensions.
@@
The estimation of causal treatment effects from data is inherently challenging due to the fundamental problem of causal inference, namely, the absence of counterfactual outcomes (Holland, 1986). In the case of observational data, this problem is exacerbated by the possible non-random assignment of treatments and the resulting confounding bias (Pearl, 2009). A real-life example is the overestimation of workplace wellness programs’ impact on employee wellbeing, where non-random incentives cause treated populations to differ from the general population (Jones et al., 2019). A naive approach may overfit on the self-selected group, missing the true causal effect. This highlights how confounding can lead to issues like Simpson’s paradox, where aggregated data misrepresents trends compared to stratified data (Jones et al., 2019). While randomized controlled trials (RCTs) are the gold standard for mitigating such biases (Verstraete et al., 2023), they are often impractical in real-world business settings because of cost, ethical, or operational reasons. Moreover, in the context of continuous treatments, with virtually infinite possible doses, setting up an RCT for dose-response estimation becomes significantly more complex (Holland-Letz and Kopp-Schneider, 2015). In contrast, observational data is often cheap and readily accessible. Therefore, methods that aim to balance the treatment and control groups have been proposed, including techniques like propensity score matching (Rosenbaum and Rubin, 1983) or covariate balancing (Johansson et al., 2016; Shalit et al., 2017).
@@
In recent years, the consideration of continuous treatments—and consequently, CADR estimation—has gained traction because of its relevance in applications where understanding the effects of varying dose levels is critical (see Table 1). Estimating CADRs is inherently more challenging than CATEs because it involves a spectrum of treatment doses, rather than values 00 and $1$ (Schwab et al., 2020). Also, with the theoretically infinite number of possible dose values, data sparsity and the non-uniformity of dose assignment could be challenging, making accurate effect estimation at each dose level difficult (Bockel-Rickermann et al., 2024). The Conditional Average Dose Effect (CADE), which can be directly derived from the CADR, generalizes the CATE for continuous treatments. Our paper specifically focuses on UM with this type of treatment. Traditionally, dose effects have been derived from average dose-response curves using methods like the Hirano-Imbens estimator (Hirano and Imbens, 2004), which extends propensity scores to continuous interventions through generalized propensity scores. However, these approaches often overlook individual-level heterogeneity in responses. Recent advancements in ML have introduced techniques for learning individualized dose-response curves (Bica et al., 2020; Schwab et al., 2020; Nie et al., 2021; Zhang et al., 2022; Bockel-Rickermann et al., 2023). However, analogously to the case of CATE estimation, these methods primarily focus on the inference step and do not incorporate an optimization step where the learned dose-response relationships are used to prescribe decisions (Fernández-Loría and Provost, 2022).
In this paper, we do not make any assumptions about the shape of the CADR. Instead, we adopt a flexible, data-driven approach. In contrast, existing work, such as (Zhou et al., 2023), relies on a strong structural assumption — the law of diminishing marginal utility — which implies that the marginal benefit of increasing the dose is strictly decreasing. While this assumption may hold in some applications, it limits model flexibility and generalizability. Our framework does not rely on such assumptions: the estimated dose-response relationships in our approach are non-parametric and can capture arbitrary, potentially non-monotonic shapes.
@@ 2.3.1 Cumulative uplift & total benefit as driver
The allocation task aims to optimize an objective value driven by (cumulative) uplift. While uplift applies to a single entity, cumulative uplift reflects the policy’s overall impact. For a policy treating $K$ entities, cumulative uplift is the sum of CATEs for discrete treatments or CADEs for continuous treatments across these treated entities. In certain applications, cumulative uplift directly aligns with the target objective, as seen in contexts where maximizing cumulative uplift is the explicit goal (Devriendt et al., 2018). However, cumulative uplift itself does not necessarily represent the objective. For example, in value-driven analytics, where uplift might disregard expected profit, and in cases where each entity may be associated with an individual benefit and treatment cost (Gubela et al., 2020; Lemmens and Gupta, 2020; Gubela and Lessmann, 2021; Verbeke et al., 2023; Haupt and Lessmann, 2022). Hence, when cumulative uplift is not the end goal in itself, allocations with lower cumulative uplift may yield a greater increase in objective value.
@@ 2.3.2 Predict-and-optimize
Although DFL approaches for UM exist (Betlei et al., 2021; Vanderschueren et al., 2024), we adopt the predict-then-optimize framework due to its simplicity, flexibility, stability, and generalizability to other treatment types. It allows using various predictive models without being restricted by a specific optimization setup, which is especially useful for complex or pre-trained models without access to the training process (Barocas et al., 2023). Moreover, it enables flexible adjustments to the objective function and integration of additional constraints other than budget limits, such as fairness considerations, allowing for different allocations under varying budget conditions. Our approach requires modifying only the optimization step, leaving the prediction model unchanged and does not require a complete model retraining.
@@ 2.3.4 Flexible optimization constraints and objectives
Balancing utility — such as profit and accuracy — against fairness, like equal treatment across demographics, remains a key challenge. Kleinberg et al. (2016) and Corbett-Davies et al. (2017) show that enhancing fairness often trades off with utility and that multiple fairness definitions may even conflict. Traditional fairness assessments emphasize output-based metrics, such as demographic parity or equal opportunity, which measure group-level outcome disparities (Barocas et al., 2023). Recent approaches consider the entire distribution of predictions or decisions, incorporating statistical moments and exploring distributional fairness (Khan et al., 2023; Han et al., 2023). In decision-making, fairness extends beyond prediction to both allocation fairness—ensuring treatment levels are independent of sensitive attributes like race or gender—and outcome fairness, which ensures equitable results across groups. Although outcome predictions may be imperfect, enforcing fairness at the level of expected outcomes has been shown to be an effective and ethically justified approach in decision-making settings (Scantamburlo et al., 2024; Corbett-Davies et al., 2017). While not ideal, these estimates still represent the best available proxy for assessing how different groups will benefit from the allocation policy. Fairness notions, which may conflict, can be incorporated as hard constraints during optimization (Corbett-Davies et al., 2017), similar to this work’s approach, or as soft constraints in multi-objective learning (Barocas et al., 2023). Recent research, including Frauen et al. (2023), has largely focused on fairness in observational data with binary sensitive attributes, while Nabi et al. (2019) apply causal inference to multi-stage decision-making under fairness constraints. However, these studies do not address fairness in the context of continuous treatments, a critical gap our research aims to fill.
@@
In managerial decision-making, profit maximization or cost reduction is often the primary goal. Cost-sensitive or value-driven methods are crucial as they incorporate asymmetric costs and benefits, aligning decisions with these business objectives (Höppner et al., 2022). Cost-sensitive approaches seek to balance costs and benefits, an aspect often ignored by standard methods (Gubela et al., 2020; Verbeke et al., 2023). In this work, costs and benefits are deterministic while assuming stochastic treatment effects, setting it apart from studies where both costs and benefits are modeled as stochastic (Verbraken et al., 2012; Verbeke et al., 2012). Cost-sensitive methods find applications across various domains, including credit risk (Bahnsen et al., 2014), fraud detection (Vasquez et al., 2022; De Vos et al., 2023), customer churn (Verbraken et al., 2012), business failure prediction (De Bock et al., 2020), and machine maintenance (Vanderschueren et al., 2023b).
@@ 3.2 Prediction step
During the prediction step, our goal is to train a model that can accurately estimate $\hat{y}(s)$, providing unbiased estimates of $\mu(s,\mathbf{x})$ and $\tau_{s}(\mathbf{x})$ across the entire domain of $S\in[0,1]$. The estimated CADR for an instance with features $\mathbf{x}$ is defined as $\hat{\mu}(s,\mathbf{x})=\mathbb{E}\left[\hat{Y}(s)\mid\mathbf{X}=\mathbf{x}\right]$. The estimated CADE for a given dose $s$ is defined as $\hat{\tau}_{s}(\mathbf{x})=\hat{\mu}(s,\mathbf{x})-\hat{\mu}(0,\mathbf{x})=\mathbb{E}\left[\hat{Y}(s)-\hat{Y}(0)\mid\mathbf{X}=\mathbf{x}\right]$. To use the CADE estimates as input for the optimization step, we discretize the CADRs into $\delta$ bins. For a given $\delta$, we define $D=\left\{\left.\frac{d-1}{\delta}\right\rvert\,d=1,\ldots,(\delta+1)\right\}$ so that the vector $\hat{\bm{\tau}}(\mathbf{x})=\Bigl{(}\hat{\tau}_{D_{d}}(\mathbf{x})\Bigr{)}_{d=1}^{\delta}$ then contains an entity’s $\delta$ CADE estimates. We discuss model training and outcome estimation in Section 4.1.